OpenStudy (anonymous):

7/8-1/4=b i think the answer is 5/8 but im not sure an i correct ?

OpenStudy (solomonzelman):

7/8 - 1/4 = b 7/8 - 2/8 = b (7-2) / 8 = b 5/8 =b You are RIGHT again :)

OpenStudy (anonymous):

thank you im never sure with fractions